hartnn (hartnn):

Can Milne-Thompson Method be used to find the Real part of f(z) when Imaginary part is given ??

hartnn (hartnn):

I know it can be used to find the imaginary part when real part is given by plugging in x=z, y=0 but what about the other way round ?

hartnn (hartnn):

If the imaginary part of the analytic function \( W=f(z)\) is \( V=x^2-y^2+x/(x^2+y^2 ) \) find the real part U

OpenStudy (kainui):

Can't you just multiply your function f(z) by i to reverse the parts around?

OpenStudy (kainui):

Suppose you have f(z)=a+bi then that means i*f(z)=-b+ai and you can give it -b to find a? Honestly I don't know I've never heard of this thing before so I can't tell you, but that's just my guess.

hartnn (hartnn):

I can try that...but I won't know for sure whether that works, as i don't have the answer....if anyone has solved such type of problem before....or anyone can give solved example from the internet, then that would be great

hartnn (hartnn):

will use your method s the last resort...thanks :)

OpenStudy (phi):

I'm not a mathematician (i.e. never heard of Milne-Thompson), but it seems reasonable for g(z)= i*f(z) that i*Re(f(z)) = Im(g(z)) and -Im(f(z))= Re(g(z))

hartnn (hartnn):

ok, thanks

OpenStudy (phi):

\[ \frac{z-1}{iz + i } \] \[ \text{z=1, } \frac{1-1}{i+ i } = 0 \\ \text{z=i, } \frac{i-1}{i^2 + i}=1 \\ \text{z=-1, } \frac{-1-1}{-i+ i }= \infty \]

hartnn (hartnn):

Put z=1 and w=0 in (A) ∴0=(a+b)/(c+d) a=-b→(B) Put z=i,w=1 in (A) 1=(ai+b)/(ci+d) ∴ai+b=ci+d a=c,b=d →(C) Put z=-1 ,w=∞ in (A) ∞=(-a+b)/(-c+d) ∴-c+d=0 c=d→(D) so, a=1,b=-1, c=d=i which means (B) and (D) are correct not (C)

hartnn (hartnn):

omg, i see a whole lot of ? in triangle boxes

OpenStudy (phi):

refresh your browser

hartnn (hartnn):

ohh! since, a,b,c,d can be complex i cannot conclude a=c, b=d from ai+b=ci+d

OpenStudy (phi):

yes, I used b= -a and d= c and solved for c in terms of a (ai - a)/(cz+c) = 1 when z= i

hartnn (hartnn):

cool, thanks :)

OpenStudy (phi):

I guess we have to remember the coefficients can be complex.

hartnn (hartnn):

yep!